OpenStudy (anonymous):

add 3/3 + 2/6 + 4/12

OpenStudy (eamier):

make all the denominator same, try 6

OpenStudy (anonymous):

Try making them all over 3 so you have 3/3+1/3+1/3 and you get 5/3

OpenStudy (anonymous):

is 5/3 the answer though

OpenStudy (eamier):

yes

OpenStudy (mertsj):

Just reduce each fraction and you have: \[1+\frac{1}{3}+\frac{1}{3}=1 \frac{2}{3}=\frac{5}{3}\]

OpenStudy (anonymous):

yep it 5/3

OpenStudy (anonymous):

alright got it thank you